# Unsupervised full-field Bayesian inference of orthotropic hyperelasticity from a single biaxial test: a myocardial case study

### v1

@@ 4 Orthotropic constitutive parameter inference
$e_{\theta}$ $=\frac{|\hat{\theta}-\theta_{\text{gt}}|}{\theta_{\text{gt}}}\qquad$ $\tilde{\theta}$ $=\frac{\hat{\theta}-\mu_{\hat{\theta}}}{\mu_{\hat{\theta}}}\qquad$ $s_{\tilde{\theta}}$ $=\frac{\sqrt{N(N-1)}}{N-2}\cdot\frac{\frac{1}{N}\sum_{n=1}^{N}(\tilde{\theta}_{n}-\mu_{\tilde{\theta}})^{3}}{\left(\frac{1}{N}\sum_{n=1}^{N}(\tilde{\theta}_{n}-\mu_{\tilde{\theta}})^{2}\right)^{3/2}}$ (28)
with $N$ equaling the amount of samples drawn from the posterior distribution. To gain insight into the shape, mean, and spread of the approximated posterior distribution, we draw 1,0001{,}000 samples from the inferred guide distribution shown in the bottom rows of Figures 6 and 7 respectively.
Figure 6: Orthotropic constitutive parameter inference from geometrically homogeneous tissue slabs with varying microstructure. From left to right: a circ-long-hom specimen, a circ-rad-hom specimen, and a rad-heli-hom specimen. From top to bottom: (i) evolution of inferred material parameters $\hat{\theta}$ over the epochs, (ii) evolution of relative parameter errors $e_{\theta}$, (iii) final relative errors at final epoch, and (iv) rescaled posterior distributions. Distributions are shown in grayscale when unbiased skewness $|s_{\theta}|<0.5$. Despite microstructural heterogeneity, geometric homogeneity limits identifiability of shear-related parameters ($a_{fs}$, $b_{fs}$), which remain poorly inferred. In contrast, isotropic and primary anisotropic parameters are consistently identified with high accuracy, exhibiting narrow and near-Gaussian posterior distributions. Increasing microstructural heterogeneity improves convergence rates and reduces inference error for shear parameters, though challenges persist for $b_{fs}$, particularly in the rad-heli-hom specimen.
Focusing on geometrically homogeneous samples (Figure 6), we evaluated three configurations with increasing microstructural heterogeneity: fully homogeneous (circ-long-hom), symmetric transmural fiber rotation (circ-rad-hom), and asymmetric helical fiber variation (rad-heli-hom). For the circ-long-hom configuration, all parameters converged rapidly except for those associated with the fiber-sheet shear coupling parameters ($a_{fs}$, $b_{fs}$). Final relative errors were minimal for isotropic parameters ($e_{a}=4.77\times 10^{-5}$, $e_{b}=-2.30\times 10^{-6}$) and remained low for anisotropic terms in $f$ and $n$ directions (i.e., $e_{a_{f}}=-5.50\times 10^{-5}$, $e_{b_{f}}=-1.47\times 10^{-6}$, $e_{a_{n}}=9.66\times 10^{-5}$ , $e_{b_{n}}=1.29\times 10^{-5}$) and the bulk modulus ($e_{k}=2.423\times 10^{-6}$). In contrast, shear-related parameters exhibited substantial errors ($e_{a_{fs}}=-141.7$, $e_{b_{fs}}=0.92$). These trends stabilized within 8,0008{,}000 epochs, except for $b_{fs}$, which remained poorly identified. Introducing symmetric transmural fiber heterogeneity, i.e. in the circ-rad-hom slab, markedly improved fiber-sheet shear parameter identification, reducing relative errors to $e_{a_{fs}}=8.05\times 10^{-3}$ and $e_{b_{fs}}=0.17$, while maintaining accuracy for isotropic and anisotropic terms (see Table 4). Convergence for $b_{fs}$ was slower, requiring up to 30,00030{,}000 epochs. Further increasing fiber and sheet deformation heterogeneity (Figure 4), that is in the rad-heli-hom slab, yielded similar improvements , though inferring $b_{fs}$ remained challenging, with a relative error of $0.374$ and no clear stabilization after 100,000100{,}000 epochs. Other parameters exhibited relative errors below $10^{-4}$, comparable to the circ-rad-hom case (Table 4). Posterior distributions (Figure 6 - bottom row) confirmed these trends. Well-identified parameters displayed narrow, symmetric distributions, with standard deviations several orders of magnitude smaller than their means. In contrast, $a_{fs}$ and $b_{fs}$ exhibited skewed or truncated distributions, particularly in the circ-long-hom configuration, indicating limited identifiability. Notably, skewness magnitudes exceeded $|s_{\theta}|>0.5$ only for $a_{fs}$ and $b_{fs}$, while other inferred parameter distributions remained near-Gaussian. Overall, these results highlight that geometric homogeneity severely limits identifiability of shear-related parameters, while even moderate microstructural heterogeneity substantially improves inference quality.
@@
Figure 7: Orthotropic constitutive parameter inference from geometrically heterogeneous tissue slabs with varying microstructure. From left to right: a circ-long-het specimen, a circ-rad-het specimen, and a rad-heli-het specimen. From top to bottom: (i) evolution of inferred material parameters $\hat{\theta}$ over the epochs, (ii) evolution of relative parameter errors $e_{\theta}$, (iii) final relative errors at final epoch, and (iv) rescaled posterior distributions. Distributions are shown in grayscale when unbiased skewness $|s_{\theta}|<0.5$. Geometric heterogeneity significantly enhances identifiability of shear-related parameters, especially $b_{fs}$, reducing relative errors to sub-percent levels in circ-rad-het and rad-heli-het specimens. Posterior distributions for these parameters become narrower and less skewed compared to homogeneous counterparts. All other parameters maintain high accuracy and exhibit symmetric, low-variance distributions. These results underline the critical role of geometric heterogeneity in resolving parameters sensitive to shear-deformation modes.
Focusing on geometrically heterogeneous samples (Figure 7), we evaluated three configurations with increasing microstructural complexity: fully homogeneous fibers (circ-long-het), symmetric transmural fiber rotation (circ-rad-het), and asymmetric helical fiber variation (rad-heli-het). For the circ-long-het configuration, all parameters were accurately identified except for $b_{fs}$ (relative error of roughly $1\%$). Relative errors remained low for isotropic parameters ($e_{a}=1.83\times 10^{-5}$, $e_{b}=1.22\times 10^{-5}$), anisotropic parameters in $f$ and $n$ directions ($e_{a_{f}}=-4.18\times 10^{-5}$, $e_{b_{f}}=9.77\times 10^{-6}$, $e_{a_{n}}=7.87\times 10^{-5}$, $e_{b_{n}}=-1.40\times 10^{-5}$), and the bulk modulus ($e_{k}=4.84\times 10^{-6}$). However, fiber-sheet shear coupling remained challenging, with errors of $e_{a_{fs}}=1.02\times 10^{-2}$ and $e_{b_{fs}}=0.879$. All parameters stabilized within 8,0008{,}000 epochs. Introducing symmetric fiber heterogeneity in the circ-rad-het slab markedly improved shear parameter inference, reducing $e_{a_{fs}}$ and $e_{b_{fs}}$ to $2.52\times 10^{-3}$ and $1.54\times 10^{-3}$, respectively. Isotropic and anisotropic parameters remained highly accurate (e.g., $e_{a}=-1.64\times 10^{-5}$, $e_{b}=-2.80\times 10^{-5}$), with convergence achieved within 5,0005{,}000 epochs. Further increasing microstructural complexity in the rad-heli-het configuration yielded comparable results. Relative errors for $a_{fs}$ and $b_{fs}$ were $1.14\times 10^{-3}$ and $6.90\times 10^{-3}$, respectively, while other parameters maintained relative errors below $10^{-4}$, as shown in the third row of Figure 7 and Table 4. Convergence required up to 10,00010{,}000 epochs, slightly slower than the circ-rad-het case. Posterior distributions (Figure 7 - bottom row) confirmed these trends. Parameters associated with isotropic and anisotropic invariants displayed narrow, near-Gaussian distributions with rescaled standard deviations between $10^{-3}$ and $10^{-5}$. In contrast, shear-related parameters ($a_{fs}$, $b_{fs}$) exhibited wider, skewed distributions, particularly in circ-long-het. However, skewness and standard deviations were notably reduced compared to homogeneous slab configurations, indicating improved identifiability. In general, these results demonstrate that geometric heterogeneity substantially enhances inference of shear-related parameters, effectively complementing microstructural variations to improve parameter identifiability.
In general, we notice in the top rows of figures 6 and 7 that the parameters $b$ , $b_{f}$, $a_{n}$ and $a_{fs}$ drop in value after roughly 1,0001{,}000 epochs, only to reach a minimal value around 5,0005{,}000 epochs and then shoot up to converge to their optimal values. Similar behavior can be seen for the remaining parameters, though they stabilize at lower values than their initial values. In the second row of these figures we can also see interesting jumps between $10$ and 5,0005{,}000 epochs. Comparing these jumps to the parameter values in the top row reveals that these are the epochs where the estimated parameter value crosses the ground-truth value without achieving a stabilizing solution. This underlines the changing posterior geometry and thus also how the loss landscape changes due to parameter optimization. The summary of all parameter fitting results under noise free conditions is provided in C.
5 Impact of measurement noise
We examined inference robustness by introducing displacement measurement noise at two levels: low ($1\times 10^{-4}\,\mathrm{mm}$) and high ($1\times 10^{-3}\,\mathrm{mm}$), applied to geometrically heterogeneous training samples, because they showed structurally better performance in the previous sections. The choice of displacement noise levels is based on similar works (Flaschel et al., 2021; Joshi et al., 2022; Thakolkaran et al., 2022; Rahmani et al., 2013). We note here that Gaussian white noise can be considered harsher than what is present in realistic data because the spatial correlation between measurements is not taken into account. Some relevant works have implemented spatially correlated noise, or adapted fictitious noise to DIC measurements directly (Jafari et al., 2025; Peyraut and Genet, 2025). By considering Gaussian white noise, we aim to validate how robust our method is.
[…]
In Figure 8, we assess parameter inference performance with low levels of displacement noise ($\sigma=10^{-4}$) for geometrically heterogeneous tissue slabs with varying levels of microstructural heterogeneity. From left to right, results are shown for (i) a fully homogeneous circ-long-het specimen, (ii) a circ-rad-het specimen with moderate transmural fiber rotation, and (iii) a rad-heli-het specimen exhibiting pronounced helical fiber variation. In the top row of the figure, we observe that the circ-long-het configuration fails to converge for the parameter value $a_{fs}$. Consequently, more stochastic behavior as well as a slight downward drift can be observed in $b_{fs}$. Although most parameter values reach good relative error values ranging from $1\times 10^{-4}$ to $5\times 10^{-3}$, the parameters associated with $I_{4fs}$ show relative errors of the same order of magnitude as the true value. This error is further highlighted by the skewness and clear truncation of the posterior distribution shown in the bottom row. The spread in values also highlights the uncertainty around these parameter values. The helically distributed fiber configurations show no such skewness in any of their parameter distributions. They do obtain higher relative errors in their parameter values. Ranging between $5\times 10^{-3}$ to $1\times 10^{-1}$ for the circ-rad configuration and $3\times 10^{-4}$ to $2\times 10^{-2}$ (excluding $a_{fs}$ and $b_{fs}$). Strikingly, the shear parameters stabilize in the circ-rad-het and rad-heli-het configurations after roughly 5,0005{,}000 and 40,00040{,}000 epochs respectively, even though their relative errors are remarkably high. Combined with the lack of skewness and truncation in their distributions, this indicates that the parameter values have moved sufficiently far away from the prior and have settled in a (sub-optimal) local minimum.
[…]
We assess the performance of the parameter fitting approach for the high-noise case in a similar way. In Figure 9, we use the same lay-out as before and observe striking differences with the low displacement noise setting. All configurations now fail to characterize the shearing parameters $a_{fs}$ and $b_{fs}$. Moreover, we now also observe that the circ-rad-het configuration fails to show stabilizing behavior for the $a_{n}$ and $b_{n}$ parameters. We also observe an increase in relative error for all parameter values in all configurations. Remarkably, the circ-long-het specimen obtains the lowest relative errors for all parameters. The circ-rad-het configuration clearly performs the worst of the three showing relative errors ranging from $2\times 10^{-1}$ to $1$. Compared to the low-noise case this indicates a sudden drop in fitting performance. All configurations show overly skew distributions in the $a_{fs}$ and $b_{fs}$ parameters, although the circ-rad configuration also shows skewness in the distributions for $a_{n}$ and $b_{n}$. The remaining distributions show standard deviations orders of magnitude lower than the mean, indicating the confidence the model has in these values. Similarly to the low-noise case we see that parameter that do stabilize, do so around 8,0008{,}000. The rad-heli-het configuration seems to show stabilizing behavior for all its parameters roughly 2,0002{,}000 epochs later.
Summarizing Figures 7, 8, and 9, we observe that the introduction of noise steadily degrades the relative error in the inferred parameter values. Furthermore, some inferred constitutive parameters start converging towards numerical zero as the amount of displacement measurement noise increases. D provides a summary of the obtained constitutive parameter values, the relative error achieved, and their standard deviation.
6 Validation

### v2

@@ 4.1 Noise-free inference: identifiability gains from microstructural and geometrical heterogeneity
$\hat{e}_{\theta}$ $=\frac{\mu_{\theta}-\theta_{\text{gt}}}{\theta_{\text{gt}}}\qquad$ $e_{\theta}$ $=\frac{\theta-\theta_{\text{gt}}}{\theta_{\text{gt}}}$ (29)
Here, we define $\mu_{\theta}=\mu^{q}$ as the mean of the underlying untruncated normal distribution used in the guide distribution. Furthermore, we define the unbiased skewness $s_{\theta}$:
@@
$s_{\tilde{\theta}}$ $=\frac{\sqrt{N(N-1)}}{N-2}\cdot\frac{\frac{1}{N}\sum_{n=1}^{N}(\tilde{\theta}_{n}-\mu_{\tilde{\theta}})^{3}}{\left(\frac{1}{N}\sum_{n=1}^{N}(\tilde{\theta}_{n}-\mu_{\tilde{\theta}})^{2}\right)^{3/2}}\qquad\text{with}$ $\tilde{\theta}$ $=\frac{\theta-\mu_{\theta}}{\mu_{\theta}}$ (30)
with $\tilde{\theta}$ denoting the relative error between samples taken from the guide distribution and the scaled posterior, $\mu_{\tilde{\theta}}$ denoting the mean of the distribution $\tilde{\theta}$, and $N$ equaling the number of samples drawn from the posterior distribution. Unless stated otherwise, we draw N=1, [^0] N=1{,}000 samples from the inferred guide distribution shown in the bottom rows of Figures 6 and 7, respectively, to visualize the shape, mean, and spread of the approximated posterior distribution. For completeness, the full set of converged parameter estimates, posterior scales (the standard deviation of the underlying and untruncated normal distribution), and relative errors for all six training specimens is reported in B (Tables 34). In the main text, we therefore focus on the parameter groups that drive the qualitative differences between Figures 6 and 7, and summarize the remaining terms compactly.
Figure 6: Noise-free parameter inference on geometrically homogeneous tissue slabs with varying microstructure. From left to right: circ-long-hom, circ-rad-hom, and rad-heli-hom specimens. From top to bottom: (i) evolution of inferred material parameters $\theta$ over optimization epochs, (ii) evolution of absolute relative location parameter errors $|\hat{e}_{\theta}|$, (iii) absolute relative errors $|e_{\theta}|$ at the final epoch with corresponding 95% credibility interval, and (iv) rescaled posterior distributions $\tilde{\theta}$ (drawn from the variational guide). The plots show rapid convergence and small final errors for most parameters across all three specimens, while the shear-coupling parameters ($a_{fs}$, $b_{fs}$) display slower convergence and broader posteriors in several cases. Distributions are shown in grayscale when the absolute unbiased skewness satisfies $|s_{\tilde{\theta}}|<0.5$.
Focusing on the geometrically homogeneous specimens (Figure 6), the plots show rapid convergence and narrow posteriors for the isotropic $(a,b,K)$ and primary anisotropic $(a_{f},b_{f},a_{n},b_{n})$ parameter groups across all three microstructural settings. Quantitatively, these parameters reach small final deviations, with the largest magnitudes remaining below $2.3\times 10^{-4}$ in this homogeneous-geometry group (Table 4). In contrast, the fiber–sheet shear-coupling terms $(a_{fs},\,b_{fs})$ are the dominant outliers. In the fully homogeneous circ-long-hom specimen, the shear-coupling terms remain poorly identified, with large final errors ($\hat{e}_{a_{fs}}=-1.42\times 10^{2}$ and $\hat{e}_{b_{fs}}=9.20\times 10^{-1}$; Table 4). Introducing transmural and helical fiber variation improves identifiability of $a_{fs}$ substantially (from $\mathcal{O}(10^{2})$ to $\mathcal{O}(10^{-3}\!-\!10^{-2})$; Table 4), while $b_{fs}$ remains the least accurate inferred parameter in the homogeneous-geometry group (e.g., $\hat{e}_{b_{fs}}=1.72\times 10^{-1}$ for circ-rad-hom and $3.74\times 10^{-1}$ for rad-heli-hom). Notably, $b_{fs}$ also converges more slowly than the remaining parameters: it stabilizes only after $\sim 3\times 10^{4}$ epochs in circ-rad-hom, and it shows no clear plateau by $10^{5}$ epochs in rad-heli-hom, whereas most other parameters stabilize within $\sim 8\times 10^{3}$ epochs (rows (i)–(ii)). These trends are consistent with the broader and more skewed posteriors for $(a_{fs},\,b_{fs})$ in Figure 6 (bottom row).
@@
Figure 7: Noise-free parameter inference on geometrically heterogeneous tissue slabs with varying microstructure. From left to right: circ-long-het, circ-rad-het, and rad-heli-het specimens. From top to bottom: (i) evolution of inferred material parameters $\theta$ over optimization epochs, (ii) evolution of absolute relative location parameter errors $|\hat{e}_{\theta}|$, (iii) absolute relative errors $|e_{\theta}|$ at the final epoch with corresponding 95% credibility interval, and (iv) rescaled posterior distributions $\tilde{\theta}$ (drawn from the variational guide). The plots show stable convergence for all parameter groups across the three specimens, and the posterior samples indicate narrower spreads for the shear-coupling parameters ($a_{fs}$, $b_{fs}$) compared with the homogeneous counterparts. Distributions are shown in grayscale when the absolute unbiased skewness satisfies $|s_{\tilde{\theta}}|<0.5$.
For the geometrically heterogeneous specimens (Figure 7), the plots show stable convergence across parameter groups and generally narrower posterior spreads than in Figure 6. For the isotropic and primary anisotropic parameters, final deviations remain small overall, with the largest magnitudes staying below $5\times 10^{-4}$ across the heterogeneous-geometry group (Table 4). The most pronounced quantitative change occurs in the inferred shear-coupling parameters. With combined geometric heterogeneity and fiber rotation, both $a_{fs}$ and $b_{fs}$ converge to small final deviations: in circ-rad-het, $\hat{e}_{a_{fs}}=-2.52\times 10^{-3}$ and $\hat{e}_{b_{fs}}=-1.54\times 10^{-3}$, and in rad-heli-het, $\hat{e}_{a_{fs}}=-1.14\times 10^{-3}$ and $\hat{e}_{b_{fs}}=6.90\times 10^{-3}$ (Table 4). In contrast, without fiber rotation (circ-long-het), $b_{fs}$ remains comparatively inaccurate ($\hat{e}_{b_{fs}}=0.879$), consistent with the broader posterior in the corresponding column of Figure 7. Compared with the homogeneous-geometry cases in Figure 6, the rotated-fiber heterogeneous specimens (circ-rad-het and rad-heli-het) reach a stable $b_{fs}$ trajectory with small final deviations within $\mathcal{O}(10^{4})$ epochs ($\sim 5\times 10^{3}$ and $\sim 10^{4}$ epochs, respectively), rather than requiring $\sim 3\times 10^{4}$ to $10^{5}$ epochs, consistent with the tighter posterior spreads in row (iv).
Across both Figures 6 and 7, several parameters exhibit non-monotonic transients before converging. In particular, multiple traces (including $b$, $b_{f}$, $a_{n}$, and $a_{fs}$) show a decrease after roughly $10^{3}$ epochs, followed by recovery towards their converged values by $\mathcal{O}(10^{4})$ epochs. Consistent excursions appear in the error plots over the same epoch range, reflecting epochs where the inferred trajectories pass through the ground-truth value (i.e., where the signed relative deviation changes sign) without immediately stabilizing. This behavior is consistent with strong parameter coupling and a nonconvex loss landscape in nonlinear orthotropic inference, where the optimizer can traverse multiple basins before stabilizing.
4.2 Robustness to displacement noise
We next assess robustness to displacement measurement noise by perturbing the measured full-field kinematics with additive Gaussian white noise at two levels: low ($\sigma_{u}=1\times 10^{-4}\,\mathrm{mm}$) and high ($\sigma_{u}=1\times 10^{-3}\,\mathrm{mm}$). We intentionally avoid modeling spatially correlated noise here. Instead, we use spatially uncorrelated Gaussian white noise as a conservative stress test of the inference framework, with noise magnitudes chosen in line with prior inverse and Bayesian identification studies [30, 31, 32, 72]. Realistic displacement fields, for example those obtained with DIC, typically exhibit spatial correlations. Related work has incorporated correlated noise models or DIC-aware fictitious noise to mimic these effects more closely [29, 73]. In this study, we deliberately adopt the simpler white-noise approximation to probe robustness under coarse, worst-case perturbations, and we apply it to the geometrically heterogeneous training specimens, since these provided the most reliable baseline inference performance in the noise-free setting (Section 4.1).
Figure 8: Parameter inference on geometrically and microstructurally heterogeneous tissue slabs under low displacement measurement noise. From left to right: circ-long-het, circ-rad-het, and rad-heli-het specimens. From top to bottom: (i) evolution of inferred material parameters $\theta$ over optimization epochs, (ii) evolution of absolute relative location parameter errors $|\hat{e}_{\theta}|$, (iii) absolute relative errors $|e_{\theta}|$ at the final epoch with corresponding 95% credibility interval, and (iv) rescaled posterior distributions $\tilde{\theta}$ (drawn from the variational guide). The plots show that most parameters maintain stable convergence under low noise, while $a_{fs}$ and $b_{fs}$ exhibit the largest variability across specimens, reflected by slower convergence and broader posteriors. Here, uncorrelated Gaussian displacement noise with standard deviation $\sigma_{u}=10^{-4}\,\mathrm{mm}$ is added to the full-field kinematics. Distributions are shown in grayscale when the absolute unbiased skewness satisfies $|s_{\tilde{\theta}}|<0.5$.
Figure 8 summarizes inference under low displacement noise for the three geometrically heterogeneous training specimens. Across all three microstructural settings, the trajectories in rows (i) and (ii) indicate that the low-noise setting largely preserves the convergence patterns of the noise-free heterogeneous baseline in Figure 7, and differences emerge most clearly in the shear-coupling terms. For circ-long-het, final relative errors for isotropic and primary anisotropic parameters stay within a narrow band, with representative values between $10^{-4}$ and $5\times 10^{-3}$ (Table 7). In contrast, the parameter trajectory for $a_{fs}$ does not settle to a clear plateau in row (i), and $b_{fs}$ exhibits increased stochasticity together with a mild downward drift. The corresponding relative errors for the shear-coupling terms are of order one, with $|\hat{e}_{a_{fs}}|\approx 1$ and $|\hat{e}_{b_{fs}}|\approx 0.9$. This behavior is mirrored in the posterior samples in row (iv), which show pronounced skewness and visible truncation for $(a_{fs},b_{fs})$, consistent with elevated uncertainty and weak practical identifiability under this noise level. Whereas in the noise-free heterogeneous setting (Figure 7) the shear parameter $a_{fs}$ could still be identified up to a relative error on the order of $10^{-2}$, the low-noise case now produces errors for both $a_{fs}$ and $b_{fs}$ of the same order of magnitude as their true values. In the helically distributed fiber specimens circ-rad-het and rad-heli-het, the posterior samples in row (iv) do not exhibit comparable skewness or truncation for the inferred parameters. Compared with the noise-free case, the final relative errors for isotropic and primary anisotropic parameters are slightly higher, remaining mostly between $5\times 10^{-3}$ and $10^{-1}$ for circ-rad-het and between $3\times 10^{-4}$ and $2\times 10^{-2}$ for rad-heli-het (excluding $a_{fs}$ and $b_{fs}$), while the majority of parameters still converge to clearly defined plateaus in rows (i) and (ii). For the shear-coupling terms, the posterior shapes of $(a_{fs},b_{fs})$ become more regular in row (iv), yet their error traces in row (ii) plateau at comparatively elevated levels. In the noise-free case (Figure 7), we achieved rotated-fiber shear-coupling errors $|\hat{e}_{a_{fs}}|$ of order $10^{-3}$ and $|\hat{e}_{b_{fs}}|$ of at most $10^{-2}$ in the circ-rad-het and rad-heli-het specimens (Table 4). Under low noise here, these errors increase to magnitudes between roughly $5\times 10^{-1}$ and order one, with the largest outlier observed for $b_{fs}$ in rad-heli-het (Table 7). Strikingly, the shear-coupling parameters stabilize in the circ-rad-het and rad-heli-het configurations after roughly 5,0005{,}000 and 40,00040{,}000 epochs respectively, even though their relative errors remain high. Combined with the lack of skewness and truncation in their posteriors, this pattern is consistent with the parameters having moved away from the prior and having settled in a suboptimal local minimum under low-noise conditions.
Figure 9: Parameter inference on geometrically and microstructurally heterogeneous tissue slabs under high displacement measurement noise. From left to right: circ-long-het, circ-rad-het, and rad-heli-het specimens. From top to bottom: (i) evolution of inferred material parameters $\theta$ over optimization epochs, (ii) evolution of absolute relative location parameter errors $|\hat{e}_{\theta}|$, (iii) absolute relative errors $|e_{\theta}|$ at the final epoch with corresponding 95% credibility interval, and (iv) rescaled posterior distributions $\tilde{\theta}$ (drawn from the variational guide). The plots show increased variability and slower stabilization for several parameters compared with the low-noise case, with the shear-coupling parameters ($a_{fs}$, $b_{fs}$) exhibiting the largest spread in the posterior samples across specimens. Here, displacement noise with standard deviation $\sigma=10^{-3}\,\mathrm{mm}$ is added to the full-field kinematics. Distributions are shown in grayscale when the absolute unbiased skewness satisfies $|s_{\tilde{\theta}}|<0.5$.
Figure 9 assesses parameter inference under high displacement measurement noise ($\sigma_{u}=10^{-3}\,\mathrm{mm}$) for the three geometrically heterogeneous specimens, using the same layout as in Figure 8. Relative to the low-noise case, the trajectories in rows (i) and (ii) show a marked loss of stability and accuracy, most prominently in the shear-coupling parameters $(a_{fs},b_{fs})$. Across all three specimens, these shear terms no longer attain reliable plateaus, their estimates for $a_{fs}$ collapse towards numerical near-zero values, and the corresponding relative errors remain of order one (Table 7). In addition, the primary anisotropic sheet-direction parameters $(a_{n},b_{n})$ in the circ-rad-het specimen fail to settle to stable values and reach relative errors close to unity, indicating a breakdown of identifiability in this parameter group under high noise. By contrast, the isotropic and primary anisotropic parameters remain comparatively more robust in circ-long-het, where final relative errors for these parameters mostly lie between $10^{-3}$ and $5\times 10^{-2}$. In the rad-heli-het specimen, these relative parameter errors fall between $5\times 10^{-2}$ and $10^{-1}$ (Table 7). The circ-rad-het configuration clearly performs worst overall, with many isotropic and primary anisotropic parameters exhibiting relative errors in the range $2\times 10^{-1}$ to $1$, representing a sharp degradation compared with both the noise-free and low-noise cases. Similar to the low-noise setting, parameters that do stabilize under high noise typically do so around 8,0008{,}000 epochs for circ-long-het and circ-rad-het, whereas the rad-heli-het specimen requires up to about 10,00010{,}000 epochs for its stable parameter groups. C provides the corresponding noisy hyperparameter and constitutive parameter estimates, together with their relative errors and posterior scale parameters (Tables 57), which quantitatively underpin the trends observed in Figures 79.
Comparing Figures 7, 8, and 9 show a clear, parameter-selective degradation of identifiability as displacement noise increases from zero to low to high levels. In the noise-free heterogeneous baseline (Figure 7), isotropic and primary anisotropic parameters achieve small final deviations, with $|\hat{e}_{\theta}|$ typically below $5\times 10^{-4}$, while the shear-coupling terms $(a_{fs},b_{fs})$ are already more sensitive in circ-long-het but remain well identified in circ-rad-het and rad-heli-het (Table 4). Under low noise (Figure 8), these isotropic and primary anisotropic errors increase by roughly one to two orders of magnitude yet mostly remain below $10^{-1}$, whereas the shear-coupling errors in circ-long-het, circ-rad-het, and rad-heli-het rise to order one or larger, consistent with the broadened and, in some cases, truncated posteriors for $(a_{fs},b_{fs})$. At high noise (Figure 9), this trend intensifies: the shear-coupling parameters lose practical identifiability in all three specimens, with $|\hat{e}_{a_{fs}}|$ and $|\hat{e}_{b_{fs}}|$ remaining of order one and the corresponding posterior samples strongly skewed, while selected primary anisotropic parameters in circ-rad-het (notably $a_{f}$, $a_{n}$, and $b_{n}$) also exhibit near-unity relative errors and non-stabilizing trajectories. By contrast, most isotropic parameters and the bulk modulus $K$ retain comparatively moderate deviations even at the highest noise level, particularly in circ-long-het and rad-heli-het, indicating a degree of robustness in the volumetric and baseline stiffness sector despite the pronounced degradation in shear-related and directional anisotropic terms.
5 Validation beyond training
We now assess how well parameter sets inferred from single-shot heterogeneous biaxial tests generalize beyond the specific training protocol. To this end, we first compare Cauchy stress responses of inferred parameter sets under classical homogeneous uni-modal deformation modes against the ground-truth parameterized Holzapfel–Ogden law. Subsequently, we probe more complex mixed-modal deformation states, sampled from a physiologically constrained invariant space, and compare the resulting strain energy between inferred and ground-truth parameters to assess generalization under heterogeneous load paths.
5.1 Uni-modal deformation mode validation